OpenStudy (anonymous):

ABCD is a parallelogram. If m angle BCD=(6x-20) and m angle DAB=(2x+80) what is the value of x A 8.3 B 12.5 C 15 D 25 I think its B help me

563blackghost (563blackghost):

Since the angles are opposites that would mean that they are both equal so our equation will be set up as.... \(\Huge{(6x-20)=(2x+80)}\) So you would simplify...

OpenStudy (anonymous):

-20+6x=2x+80 4x= 100 +2x +2x. 4. 4. X=25 did Ido it right @563blackghost

563blackghost (563blackghost):

Correct it is 25 ^^

OpenStudy (anonymous):

Thank u 😁 @563blackghost

563blackghost (563blackghost):

np ^^

563blackghost (563blackghost):

\(\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\\\color{white}{.}\\\Huge\sf\color{blue}{~~~~Welcome~to~OpenStudy!~\ddot\smile}\\\color{white}{.}\\\\\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\) \(\Large{By:thomaster}\) Welcome to Openstudy! And I hope you will enjoy this helpful site! ^^ Please check out the code of conduct to learn the rules of this site xD http://openstudy.com/code-of-conduct